OpenStudy (anonymous):

find all values of x, 0

OpenStudy (anonymous):

your real job is to solve \[\cos(x)-\cot(x)\geq0\] or \[\cos(x)\geq \cot(x)\]

OpenStudy (anonymous):

..?

OpenStudy (anonymous):

well apparently this is true on the interval \[[\frac{\pi}{2},\pi]\]

OpenStudy (anonymous):

here they are graphed together. you can see where cosine is larger than cotangent http://www.wolframalpha.com/input/?i=y%3Dcos%28x%29%2Cy%3Dcot%28x%29

OpenStudy (anonymous):

actually, if we want to cheat we might as well do it right. here is the answer, just be careful that you can understand what it is telling you http://www.wolframalpha.com/input/?i=cos%28x%29-cot%28x%29%3E0

OpenStudy (anonymous):

thankk youu haha